Tech Meeting Minutes 20110616
Present: Peter, Sam, Stuart, Mark
Agenda Location: http://indico.cern.ch/conferenceDisplay.py?confId=138124
|Table of contents|
Gold Star to Edinburgh for no tickets.
Gold Star to Durham for getting Cream Installed.
Silver Star to Glasgow.
CREAM CE installed. Testing on going.
SL4 to SL5 update ongoing
Computing Futures Conference running so no weekly update from
Remained Operational during power work
Svr016 issues ongoing. Good progress made stabilising system until replacement hardware delivered.
Pheno Tickets ongoing. Issue appears to be submission into SARA by user but still under investigation.
Iperf Test with FZK and Manchester ongoing. Glasgow throughput to FZK significantly higher than Manchester.
ARC being installed on a "real" server.
EMI Cream release ongoing.
Mark will be in CERN on the 23rd (Unfortunately Lembit Opik was unavailable for the 23rd. Therefore the Data Storage Manager Dr Sam Skipsey will host.)